OpenStudy (anonymous):

f(x)=x+5, g(x)=-4x^2+11x+2 Find (fg)(3)

OpenStudy (anonymous):

f(3) = 8 g(3) = -1 (fg)(3) = 8 x -1 = -8

OpenStudy (anonymous):

where did you get the (fg)(3)=8x-1=-8

OpenStudy (phi):

moneybird is saying (fg)(3) means f(3) g(3), the x means times. Does (fg) mean multiply, or composition?

OpenStudy (anonymous):

multiply

OpenStudy (phi):

So the answer is 8 times -1 = -8 you could also multiply out (x+5)(-4x^2+11x+2), and then plug in x=3, and you would get -8. However, that's a lot of work.